# Unlocking Chemical Insights: Superior Molecular Representations from Intermediate Encoder Layers

### v1

@@ 4.2 Superiority of Intermediate Layers Persists with Finetuning
Next, we investigated whether the observed benefits of intermediate layers translate to a scenario where the encoder, up to the selected layer, and a task-specific prediction head are jointly finetuned. Figure 2 shows the relative difference in the test metric when finetuning to the optimal intermediate layer compared to finetuning to the final-layer, again with each cell representing a specific model-task pair and showing the numerical percentage gain. Layer-by-layer performance visualizations for all models and tasks are provided in Appendix C.
@@
Figure 2: Percentage change in test metric achieved by finetuning up to the best performing intermediate layer compared to finetuning up to the final-layer. A "–" symbol indicates that the final-layer was better.
@@
We observed pronounced benefits from employing intermediate layers even in this finetuning context. On average, choosing the optimal intermediate layer led to an 8.5% improvement across all model-task combinations, with approximately 71% of pairs benefiting from this approach. The gains were model-dependent but consistently substantial: Pos-EGNN saw the highest average improvement at 11.2%, closely followed by Uni-Mol at 9.9%, while MolFormer and OrbV3-Direct each experienced average gains of 8.4%. OrbV3-Conservative exhibited smaller, yet still notable, gains of 5.8%.
Tasks evaluated with Spearman correlation benefited the most from intermediate layers. Notably, the half-life-obach task improved by 33.8% on average, and the vdss-lombardo task by 22%, underscoring substantial advantages in tasks sensitive to rank-order predictions. In contrast, certain tasks exhibited limited improvement from intermediate layers: cyp3a4-veith improved only marginally by 0.9%, and bbb-martins increased by just 2%. Consistent with the variability observed in our fixed embedding analysis, the finetuning results also diverged from findings in other domains where there is an optimal depth for downstream performance.
@@
Remarkably, several intermediate layers deliver better results than previously reported state-of-the-art benchmarks even without extensive hyperparameter tuning. For instance, in the lipophilicity-astrazeneca task, Pos-EGNN’s layers 3 and 4 achieve an MAE of 0.459 versus the SOTA of 0.467 doi:10.1021/acs.jcim.9b00237 . In the ppbr-az task, MolFormer’s layer 8 attains an MAE of 7.221 compared to the SOTA of 7.526 notwell2023admetpropertypredictioncombinations . For cyp3a4-substrate-carbonmangels, MolFormer’s layer 9 reaches 0.699 AUROC and Uni-Mol’s layer 13 attains 0.688, both exceeding the previously reported 0.662 10.1093/bioinformatics/btaa1005 . Finally, on cyp2c9-substrate-carbonmangels, Pos-EGNN’s layer 3 records an AUCPR of 0.460 versus 0.441 Turon2022.12.13.520154 .
@@
Beyond predictive improvements, leveraging intermediate layers substantially reduces the number of trainable parameters. Because the number of parameters scales nearly linearly with encoder depth, stopping finetuning at an intermediate layer proportionally reduces the computational burden, as all layers above the selected cutoff are pruned from the network and excluded from both forward and backward passes. This yields models that not only train faster but also require fewer computational resources, facilitating efficient experimentation and deployment without compromising, and frequently even enhancing, predictive performance.
4.3 Fixed Embedding Performance as a Proxy for Finetuning Layer Selection
@@
For each molecular encoder-task pair in the TDC benchmark, we computed the Pearson correlation between the layer-wise fixed embedding scores and the corresponding finetuned scores, thereby quantifying how well embedding performance anticipates finetuning gains. Figure 3 illustrates how strongly fixed embedding performance tracks finetuned performance. The left panel shows one representative scatter plot for MolFormer on the cyp2c9-veith task, where each data point corresponds to a layer, with a Pearson correlation of 0.87. This near linear relationship indicates that layers scoring highest in fixed embedding metric are typically the same layers that achieve the top metric after full finetuning. Individual scatter plots for all model-task combinations are provided in Appendix D – H.
The right panel aggregates this analysis across all 110 model–task pairs. The distribution is heavily skewed toward positive correlations: the median is 0.60, signaling a strong alignment for half of the model-task pairs. The mean correlation of 0.41 approaches the lower bound of a useful signal, indicating fixed embedding scores can serve as a practical first-order filter to prioritize layers for finetuning, substantially reducing the search space.
@@
Figure 3: Left: Example of scatter plot of fixed embedding AUCPR vs. finetuned AUCPR for each MolFormer layer on task cyp2c9-veith. Each point is annotated with its corresponding layer number. Right: Histogram of embedding-to-finetuned correlations for all 110 model-task combinations.
@@ 5 Conclusion
This work systematically challenges the common practice of relying on final-layer representations from pretrained molecular encoders. Our comprehensive study, spanning five diverse molecular encoders and 22 ADMET tasks, demonstrates that intermediate layers frequently offer superior downstream performance. This superiority is evident both with fixed embeddings, where in over 81% of model-task combinations an intermediate layer outperformed the final-layer, yielding an average 5.4% performance increase. During finetuning, intermediate layers led to notable improvements in over 71% of cases, with an average gain of 8.5%, alongside significant savings in parameters and compute time. These findings challenge the prevailing final-layer convention, highlighting substantial potential for predictive performance enhancement, while the strong positive correlation, median Pearson 0.60, supports an efficient evaluate-then-finetune strategy.
While our findings are encouraging, we acknowledge certain limitations. Although constrained by computational resources to five encoders and ADMET tasks, extending this analysis to a broader set of molecular encoders and chemical domains could further validate our findings. Future work with additional architectures, multiple random seeds, and more exhaustive hyperparameter searches would provide deeper insights and more robust generalizations.
In conclusion, our results strongly advocate for a shift beyond the final-layer default, urging exploration of the rich representational landscape within molecular encoders. By systematically evaluating intermediate layers, the field can unlock significant performance improvements and develop more efficient modeling pipelines in molecular science.
@@ Appendix A TDC-ADMET Task Summary
Appendix B Validating TabPFN Against a Strong Tree-Based Baseline
To ensure that TabPFN is an adequate surrogate for assessing embedding quality, we compared it to a state-of-the-art gradient-boosted decision-tree model. Specifically, we followed the CatBoost pipeline described in [47]: each molecule is represented by the concatenation of ECFP-1024 [51], Avalon-1024 [52], ErG-315 [53], 200 descriptors from RDKit [54], and a 300-dimensional GIN embedding [55], yielding a 2863-feature input that achieved top-3 accuracy on 19/22 TDC-ADMET leaderboards.
@@ Appendix B Validating TabPFN Against a Strong Tree-Based Baseline
Figure 4: Comparison of CatBoost and TabPFN across a suite of ADMET benchmarks. The vertical dashed line separates the MAE-based tasks on the left (lower is better) from the remaining tasks on the right (higher is better). To display ppbr-az on a 0–1 scale, its MAE was divided by 10.
@@ Appendix C Layer-wise Performance
Each figure in this appendix presents the complete layer-wise trajectories for all five encoders on a given TDC task. The left panel reports the fixed embedding performance obtained by freezing the encoder at successive depths and training a lightweight TabPFN surrogate on top; the right panel shows the outcome when the encoder is finetuned up to the same depth together with a task-specific prediction head. Finetuning hyperparameters for each model and layer were determined by conducting a limited grid search over a few learning rates and hidden layer sizes. Depth is expressed as a percentage of the total number of encoder blocks (0% = first encoder layer; 100 % = final block), and performance is plotted according to its task metric.
@@
Comparing the two panels reveals that fixed embedding curves frequently, but not always, anticipate the relative ordering observed after fine-tuning. In some tasks (e.g. lipophilicity-astrazeneca and ppbr-az) the layer that minimizes/maximizes the given metric in the surrogate setting also yields the best or near best score after finetuning; in others, the correspondence weakens, suggesting that perhaps a broader hyperparameter search is needed. Additionally, it should be noted that for a small percentage of cases, such as vdss-lombardo and half-life-obach, finetuning provided lower performance than using fixed embeddings. This variability likely contributes to the long tail observed in the histogram of Pearson correlations (Figure 3).
@@
Finally, no single region of the network consistently outperforms the rest. Optimal depths vary not only across encoders but also across tasks within the same encoder. This heterogeneity reinforces the utility of our two-step protocol: inexpensive fixed embedding evaluation first narrows the set of candidate layers, after which targeted finetuning can focus computational resources on the most promising depths.
@@
Figure 5: MAE (↓, the lower the better) on the lipophilicity-astrazeneca dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 6: MAE (↓, the lower the better) on the caco2-wang dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 7: MAE (↓, the lower the better) on the ld50-zhu dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 8: MAE (↓, the lower the better) on the solubility-aqsoldb dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 9: MAE (↓, the lower the better) on the ppbr-az dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 10: AUROC (↑, the higher the better) on the bbb-martins dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 11: AUROC (↑, the higher the better) on the hia-hou dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 12: AUROC (↑, the higher the better) on the pgp-broccatelli dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 13: AUROC (↑, the higher the better) on the bioavailability-ma dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 14: AUROC (↑, the higher the better) on the cyp3a4-substrate-carbonmangels dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 15: AUROC (↑, the higher the better) on the ames dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 16: AUROC (↑, the higher the better) on the herg dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 17: AUROC (↑, the higher the better) on the dili dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 18: SPEARMAN (↑, the higher the better) on the vdss-lombardo dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 19: SPEARMAN (↑, the higher the better) on the half-life-obach dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 20: SPEARMAN (↑, the higher the better) on the clearance-microsome-az dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 21: SPEARMAN (↑, the higher the better) on the clearance-hepatocyte-az dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 22: SPEARMAN (↑, the higher the better) on the cyp2d6-substrate-carbonmangels dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 23: AUCPR (↑, the higher the better) on the cyp2c9-substrate-carbonmangels dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 24: AUCPR (↑, the higher the better) on the cyp2d6-veith dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 25: AUCPR (↑, the higher the better) on the cyp3a4-veith dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
@@
Figure 26: AUCPR (↑, the higher the better) on the cyp2c9-veith dataset as a function of encoder depth (% depth) for fixed embeddings (left) versus finetuned embeddings (right).
Appendix D MolFormer: Fixed Embedding vs. Finetuned Scatter Plots
@@
This appendix section presents a series of scatter plots dedicated to the MolFormer model. Each plot illustrates the relationship between fixed embedding performance and full finetuning performance across the model’s different layers for a specific downstream task. The plots were ordered by Pearson correlation.
@@
Figure 27: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 28: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 29: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 30: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 31: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 32: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 33: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 34: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 35: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 36: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 37: Fixed embeddings vs. finetuned performance for every layer of MolFormer.
Appendix E Uni-Mol: Fixed Embedding vs. Finetuned Scatter Plots
@@
This appendix section presents a series of scatter plots dedicated to the Uni-Mol model. Each plot illustrates the relationship between fixed embedding performance and full finetuning performance across the model’s different layers for a specific downstream task. The plots were ordered by Pearson correlation.
Figure 38: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 39: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 40: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 41: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 42: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 43: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 44: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 45: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 46: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 47: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 48: Fixed embeddings vs. finetuned performance for every layer of Uni-Mol.
@@
Appendix F OrbV3-Direct: Fixed Embedding vs. Finetuned Scatter Plots
@@
This appendix section presents a series of scatter plots dedicated to the OrbV3-Direct model. Each plot illustrates the relationship between fixed embedding performance and full finetuning performance across the model’s different layers for a specific downstream task. The plots were ordered by Pearson correlation.
@@
Figure 49: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 50: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 51: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 52: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 53: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 54: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 55: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 56: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 57: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.

### v2

@@ 4.3 Superiority of Intermediate Layers Persists with Finetuning
Figure 3: Percentage change in test metric achieved by finetuning up to the best intermediate layer compared to finetuning up to the final-layer. Positive values mean the best non-final layer improves over the final-layer. Negative values mean the final-layer outperforms the best non-final layer.
Next, we investigated whether the observed benefits of intermediate layers translate to a scenario where the encoder, up to a selected layer, and a task-specific prediction head are jointly finetuned. Figure 3 shows the relative difference in the test metric when finetuning to the optimal intermediate layer compared to finetuning to the final-layer, again with each cell representing a specific model-task pair and showing the numerical percentage gain. Layer‐by‐layer performance visualizations for all models and tasks are provided in Appendix C.
@@
We observed pronounced benefits from employing intermediate layers even in this finetuning context. Averaged across all pairs, the optimal intermediate layer improves the test metric by 8.5%, with 71% of pairs showing gains. The gains were model-dependent but consistently substantial: Pos-EGNN saw the highest average improvement at 11.2%, closely followed by Uni-Mol at 9.9%, while MolFormer and OrbV3-Direct each experienced average gains of 8.4%. OrbV3-Conservative exhibited smaller, yet still notable, gains of 5.8%. Consistent with the information flow diagnostics and frozen embedding results, no single depth was best, which point to broadly similar interior representations and a specialized final block.
@@
Tasks evaluated with Spearman correlation benefited the most from intermediate layers (e.g., half-life-obach +33.8%, vdss-lombardo +22%), although negatives do occur and clearance-hepatocyte-az shows both large gains and large drops across models, underscoring task-specific sensitivity to depth. Moreover, several intermediate layers deliver better results than prior SOTA benchmarks even without extensive hyperparameter tuning. For instance, Pos-EGNN Layer 3 on lipophilicity-astrazeneca (MAE 0.459 vs. 0.467), MolFormer Layer 8 on ppbr-az (MAE 7.221 vs. 7.526) and Uni-Mol Layer 13 on cyp3a4-substrate-carbonmangels (AUROC 0.688 vs. 0.662)notwell2023admetpropertypredictioncombinations ; Turon2022.12.13.520154 ; doi:10.1021/acs.jcim.9b00237 ; 10.1093/bioinformatics/btaa1005 .
@@
Beyond predictive improvements, leveraging intermediate layers substantially reduces the number of trainable parameters. Since the number of parameters scales nearly linearly with encoder depth, stopping finetuning at an intermediate layer proportionally reduces the computational burden. For instance, MolFormer truncated at layer 8 of its 12 total layers prunes the final third of the network, decreasing trainable parameters by approximately 33%. This yields models that not only train faster but also require fewer computational resources, facilitating efficient experimentation and deployment without compromising, and frequently even enhancing, predictive performance.
4.4 Frozen Embedding Performance as a Proxy for Finetuning Layer Selection
@@
For each encoder-task pair in the TDC, we computed the Pearson correlation between the layer‑wise frozen embedding scores and the corresponding finetuned scores, thereby quantifying how well embedding performance anticipates finetuning gains. Figure 4 shows the correlation distribution. The left panel shows one representative scatter plot for MolFormer on the cyp2c9-veith task, where each data point corresponds to a layer, with a Pearson correlation of 0.87. This near linear relationship indicates that layers scoring highest in frozen embedding metric are typically the same layers that achieve the top metric after finetuning. Individual scatter plots for all model-task combinations are provided in Appendix D – H. The right panel aggregates this analysis across all 110 model–task pairs. The distribution is heavily skewed toward positive correlations: the median is 0.60, signaling a strong alignment for half of the model‑task pairs and indicating frozen embedding scores can serve as a practical first-order filter to prioritize layers for finetuning, substantially reducing the search space.
@@
Figure 4: Left: Example of scatter plot of frozen embedding AUCPR vs. finetuned AUCPR for each MolFormer layer on task cyp2c9-veith. Each point is annotated with its corresponding layer number. Right: Histogram of embedding-to-finetuned correlations for all 110 model-task combinations.
@@ 5 Conclusion
This work systematically challenges the common practice of relying on final-layer representations from pretrained molecular encoders. Our comprehensive study, spanning five diverse molecular encoders and 22 ADMET tasks, demonstrates that intermediate layers frequently offer superior downstream performance. This superiority is evident when evaluating frozen embeddings, where in over 81% of model-task combinations an intermediate layer outperformed the final-layer, yielding an average 5.4% performance increase. With finetuning, intermediate layers led to notable improvements in over 71% of cases, with an average gain of 8.5%, alongside significant savings in parameters and compute time. These empirical gains are supported by our label-free analysis, which shows that the final-layer’s specialization often leads to representational compression, leaving intermediate layers with more general transferable features.
While our findings are encouraging, we acknowledge certain limitations. Although constrained by computational resources to five encoders and 22 ADMET tasks, extending this analysis to a broader set of molecular encoders and chemical domains could further validate our findings. Future work with additional architectures, multiple random seeds, and more exhaustive hyperparameter searches would provide deeper insights and more robust generalizations. Nevertheless, our results strongly advocate for a shift beyond the final-layer default, urging exploration of the rich representational landscape within molecular encoders.
@@ Appendix A TDC-ADMET Task Summary
Appendix B Validating TabPFN Against a Strong Tree‑Based Baseline
To ensure that TabPFN is an adequate surrogate for assessing embedding quality, we compared it to a state-of-the-art gradient‑boosted decision‑tree model. Specifically, we followed the CatBoost pipeline described in [50]: each molecule is represented by the concatenation of ECFP‑1024 [54], Avalon‑1024 [55], ErG‑315 [56], 200 descriptors from RDKit [57], and a 300‑dimensional GIN embedding [58], yielding a 2863‑feature input that achieved top‑3 accuracy on 19/22 TDC‑ADMET leaderboards.
@@ Appendix B Validating TabPFN Against a Strong Tree‑Based Baseline
Figure 5: Comparison of CatBoost and TabPFN across a suite of ADMET benchmarks. The vertical dashed line separates the MAE‐based tasks on the left (lower is better) from the remaining tasks on the right (higher is better). To display ppbr‐az on a 0–1 scale, its MAE was divided by 10.
@@ Appendix C Layer-wise Performance
Each figure in this appendix presents the complete layer‑wise trajectories for all five encoders on a given TDC task. The left panel reports the frozen embedding performance obtained by freezing the encoder at successive depths and training a lightweight TabPFN surrogate on top; the right panel shows the outcome when the encoder is finetuned up to the same depth together with a task‑specific prediction head. Finetuning hyperparameters for each model and layer were determined by conducting a limited grid search over a few learning rates and hidden layer sizes. Depth is expressed as a percentage of the total number of encoder blocks (0% = first encoder layer; 100 % = final block), and performance is plotted according to its task metric.
@@
Comparing the two panels reveals that frozen embedding curves frequently, but not always, anticipate the relative ordering observed after fine‑tuning. In some tasks (e.g. lipophilicity‑astrazeneca and ppbr-az) the layer that minimizes/maximizes the given metric in the surrogate setting also yields the best or near best score after finetuning; in others, the correspondence weakens, suggesting that perhaps a broader hyperparameter search is needed. Additionally, it should be noted that for a small percentage of cases, such as vdss-lombardo and half-life-obach, finetuning provided lower performance than using frozen embeddings. This variability likely contributes to the long tail observed in the histogram of Pearson correlations (Figure 4).
@@
Finally, no single region of the network consistently outperforms the rest. Optimal depths vary not only across encoders but also across tasks within the same encoder. This heterogeneity reinforces the utility of our two‑step protocol: inexpensive frozen embedding evaluation first narrows the set of candidate layers, after which targeted finetuning can focus computational resources on the most promising depths.
@@
Figure 6: MAE (↓, the lower the better) on the lipophilicity-astrazeneca dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 7: MAE (↓, the lower the better) on the caco2-wang dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 8: MAE (↓, the lower the better) on the ld50-zhu dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 9: MAE (↓, the lower the better) on the solubility-aqsoldb dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 10: MAE (↓, the lower the better) on the ppbr-az dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 11: AUROC (↑, the higher the better) on the bbb-martins dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 12: AUROC (↑, the higher the better) on the hia-hou dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 13: AUROC (↑, the higher the better) on the pgp-broccatelli dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 14: AUROC (↑, the higher the better) on the bioavailability-ma dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 15: AUROC (↑, the higher the better) on the cyp3a4-substrate-carbonmangels dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 16: AUROC (↑, the higher the better) on the ames dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 17: AUROC (↑, the higher the better) on the herg dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 18: AUROC (↑, the higher the better) on the dili dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 19: SPEARMAN (↑, the higher the better) on the vdss-lombardo dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 20: SPEARMAN (↑, the higher the better) on the half-life-obach dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 21: SPEARMAN (↑, the higher the better) on the clearance-microsome-az dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 22: SPEARMAN (↑, the higher the better) on the clearance-hepatocyte-az dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 23: SPEARMAN (↑, the higher the better) on the cyp2d6-substrate-carbonmangels dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 24: AUCPR (↑, the higher the better) on the cyp2c9-substrate-carbonmangels dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 25: AUCPR (↑, the higher the better) on the cyp2d6-veith dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 26: AUCPR (↑, the higher the better) on the cyp3a4-veith dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
@@
Figure 27: AUCPR (↑, the higher the better) on the cyp2c9-veith dataset as a function of encoder depth (% depth) for frozen embeddings (left) versus finetuned embeddings (right).
Appendix D MolFormer: Frozen Embedding vs. Finetuned Scatter Plots
@@
This appendix section presents a series of scatter plots dedicated to the MolFormer model. Each plot illustrates the relationship between frozen embedding performance and full finetuning performance across the model’s different layers for a specific downstream task. The plots were ordered by Pearson correlation.
@@
Figure 28: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 29: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 30: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 31: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 32: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 33: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 34: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 35: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 36: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 37: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
@@
Figure 38: Frozen embeddings vs. finetuned performance for every layer of MolFormer.
Appendix E Uni-Mol: Frozen embedding vs. Finetuned Scatter Plots
@@
This appendix section presents a series of scatter plots dedicated to the Uni-Mol model. Each plot illustrates the relationship between frozen embedding performance and full finetuning performance across the model’s different layers for a specific downstream task. The plots were ordered by Pearson correlation.
Figure 39: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 40: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 41: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 42: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 43: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 44: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 45: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 46: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 47: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 48: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
Figure 49: Frozen embeddings vs. finetuned performance for every layer of Uni-Mol.
@@
Appendix F OrbV3-Direct: Frozen embedding vs. Finetuned Scatter Plots
@@
This appendix section presents a series of scatter plots dedicated to the OrbV3-Direct model. Each plot illustrates the relationship between frozen embedding performance and full finetuning performance across the model’s different layers for a specific downstream task. The plots were ordered by Pearson correlation.
@@
Figure 50: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 51: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 52: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 53: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 54: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 55: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 56: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 57: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 58: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.